# Polynomial time sampling from log-smooth distributions in fixed dimension under semi-log-concavity of the forward diffusion with application to strongly dissipative distributions

### v2

@@ A.3 Proof of Proposition 7
A.4 Proof of Proposition 9
@@
In particular, we remark that $\nabla\log(Q_{t}(f))=\nabla\log(p_{t}^{V})-\nabla\log(\pi)$. Now, the quantity $\nabla\log(Q_{t}(f))$ was studied in Mikulincer and Shenfeld, [35] and they prove in Lemma 5 that for all $z$
$\nabla^{2}\log(Q_{t}(f))(z)\succeq\frac{(1-L)e^{-2t}}{(1-e^{-2t})(L-1)+1}I_{d}\,,$
@@
$\nabla^{2}\log(\pi)(z)-\nabla^{2}\log(p_{t}^{V})(z)\preceq\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}I_{d}\,.$
A.5 Proof of Lemma 10
@@
$\partial\Phi_{t}=\Phi_{t}\left(\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\right)\,.$
@@
$\partial\log(p_{t}^{V})=\Delta\log(p_{t}^{V})+\|\nabla\log(p_{t}^{V})\|^{2}-\langle\nabla\log(p_{t}^{V}),\nabla\log(\pi)\rangle-\Delta\log(\pi)\,.$
@@
$\partial\log(\Phi_{t})=\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(p_{t}^{2V})\|^{2}-\|\nabla\log(p_{t}^{V})\|^{2}\,.$
@@
$\partial\log(\Phi_{t})=\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\,.$
@@
$\partial m_{0}(\Phi_{t})$ $=\int\partial\Phi_{t}\mathrm{d}z$
@@
where we used Lemma 20 to compute $\partial\Phi_{t}$. Using integration by parts, the first term reads
@@
$\partial m_{0}(\Phi_{t})=2\int\langle\nabla\log(p_{t}^{V}),\nabla\Phi_{t}\rangle\mathrm{d}z-\int\langle\nabla\log(\pi),\nabla\Phi_{t}\rangle\mathrm{d}z\,.$
@@
$\partial m_{0}(\Phi_{t})$ $=\int\Delta\log(\pi)\Phi_{t}\mathrm{d}z-2\int\Delta\log(p_{t}^{V})\Phi_{t}\mathrm{d}z$
@@
Using Proposition 9, since $V$ is $L$-smooth, the term $(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))$ can be upper-bounded uniformly by $\frac{d(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}$ so we eventually get
$\partial m_{0}(\Phi_{t})\leq dm_{0}(\Phi_{t})\left(1+\frac{2(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}\right)\,.$
Hence we can use Gronwall lemma which yields
$m_{0}(\Phi_{t})\leq m_{0}(\Phi_{0})\exp\left(d\int_{0}^{t}\left(1+\frac{2(L-1)e^{-2s}}{(1-e^{-2s})(L-1)+1}\right)\mathrm{d}s\right)\,.$
@@
Finally, let us compute the integral above. Making the change of variable $u=e^{-2s}(L-1)$ we have $\mathrm{d}u=-2(L-1)e^{-2s}\mathrm{d}s$ which yields
$\int_{0}^{t}\frac{2(L-1)e^{-2s}}{(1-e^{-2s})(L-1)+1}\mathrm{d}s$ $=-\int_{L-1}^{(L-1)e^{-2t}}\frac{1}{L-u}\mathrm{d}u$
$=\left[\log(L-u)\right]_{L-1}^{(L-1)e^{-2t}}$
$=\log(L-(L-1)e^{-2t})$
$=\log(L(1-e^{-2t})+e^{-2t})\,.$
@@
$m_{0}(\Phi_{t})\leq\frac{e^{td}(Z_{V})^{2}}{Z_{2V}}(L(1-e^{-2t})+e^{-2t})^{d}\,.$
A.6 Upper bound on the second moment of the ratio
@@
Let $z_{t}$ be a point where $\Phi_{t}$ attains its maximum and denote $M_{t}=\log(\Phi_{t})(z_{t})$. By the implicit function theorem, $z_{t}$ is differentiable hence we can compute $\partial M_{t}$ as
$\partial M_{t}$ $=\partial\log(\Phi_{t})(z_{t})+\langle\partial z_{t},\nabla\log(\Phi_{t})(z_{t})\rangle$
@@
Since $z_{t}$ is a maximum, we have in particular $\Delta\Phi_{t}(z_{t})\leq 0$ which implies that $M_{t}$ decreases. ∎
@@
Let $V$ be a $C^{2}$ potential such that $\nabla^{2}V\preceq LI_{d}$. Denoting $\mu\propto e^{-V}$, it holds that $\frac{\mathrm{d}\mu}{\mathrm{d}z}\leq\left(\frac{\beta}{2\pi}\right)^{\frac{d}{2}}$.
@@
where $V_{*}$ the minimum of $V$ attained for some $z_{*}$. By definition $e^{-(V(z)-V_{*})}\leq 1$ for all $z$. Furthermore, since $V$ verifies $\nabla^{2}V\preceq LI_{d}$, we are ensured that
$V(z)-V_{*}\leq L\frac{\|z-z_{*}\|^{2}}{2}\,,$
@@
$\frac{1}{\int_{z}e^{-(V(z)-V_{*})}\mathrm{d}z}\leq\left(\frac{L}{2\pi}\right)^{\frac{d}{2}}\,.$
Using the previous result, we can derive an upper-bound on the integrated squared gradient at 00.
@@
Let $\mu\propto e^{-V}$ be a $L$-log smooth measure. Denoting $\mu(0)$ the density of $\mu$ with respect to the Lebesgue measure at 00, it holds that
$\int_{0}^{t}\|\nabla\log(p_{s}^{V})\|^{2}(0)\mathrm{d}s\leq-\log\left(\mu(0)\right)+\frac{d}{2}\log\left(\frac{L}{2\pi}\right)\,.$
@@
$\partial p_{t}^{V}=\nabla\cdot\left(p_{t}^{V}\nabla\log\left(\frac{p_{t}^{V}}{\pi}\right)\right)\,,$
@@
$\partial\log(p_{t}^{V})$ $=\frac{\Delta p_{t}^{V}}{p_{t}^{V}}-\langle\nabla\log(p_{t}^{V}),\nabla\log(\pi)\rangle-\Delta\log(\pi)$
@@
$\partial\log(p_{t}^{V})(0)=\Delta\log(p_{t}^{V})(0)+\|\nabla\log(p_{t}^{V})\|^{2}(0)-\Delta\log(\pi)(0)\,,$
@@
Using the uniform upper-bound of Proposition 9, the second term can be upper-bounded as
$\int_{0}^{t}\Delta\log(\pi)(0)-\Delta\log(p_{t}^{V})(0)\mathrm{d}s$ $\leq d\int_{0}^{t}\frac{(L-1)e^{-2s}}{(1-e^{-2s})(L-1)+1}\mathrm{d}s$
$=\frac{d}{2}(L-1)\int_{e^{-2t}}^{1}\frac{1}{L-u(L-1)}\mathrm{d}u$
$=\frac{d}{2}\log(L(1-e^{-2t})+e^{-2t})\,.$
Furthermore, Proposition 9 shows that $\nabla^{2}\log(p_{t}^{V})\preceq\frac{L}{L(1-e^{-2t})+e^{-2t}}$. Thus, using Proposition 22, we recover that $\log(p_{t}^{V})(0)\leq\frac{d}{2}\log\left(\frac{L}{L(1-e^{-2t})+e^{-2t}}\right)-\frac{d}{2}\log(2\pi)$. In particular, we recover
$\int_{0}^{t}\|\nabla\log(p_{s}^{V})\|^{2}(0)\mathrm{d}s\leq-\log\left(\frac{\mathrm{d}\mu}{\mathrm{d}x}(0)\right)+\frac{d}{2}\log(L)-\frac{d}{2}\log(2\pi)\,.$
@@
Let $V$ be a $C^{2}$, $L$-smooth potential such that $\mu\propto e^{-V}$ has finite second moment $m_{2}(\mu)$. It holds that
$m_{1}(\Phi_{t})\leq\frac{(Z_{V})^{2}}{Z_{2V}}e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d}\left(\sqrt{m_{2}(\mu)}+\sqrt{-2\log\left(\mu(0)\right)+d\log\left(\frac{L}{2\pi}\right)}+2\sqrt{dL}\right)\,.$
@@
$\partial m_{1}(\Phi_{t})=\int\Phi_{t}\left(\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\right)\|z\|\mathrm{d}z\,.$
@@
$\partial m_{1}(\Phi_{t})=\int\langle 2\nabla\log(p_{t}^{V})-\nabla\log(\pi),\nabla\Phi_{t}\rangle\|z\|\mathrm{d}z-\int\langle\nabla\Phi_{t},\frac{z}{\|z\|}\rangle\mathrm{d}z\,.$
@@
Using the upper bound given in Proposition 9, we get $2\int\Phi_{t}(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|\mathrm{d}z\leq 2d\frac{(L-1)e^{-2t}}{(L-1)(1-e^{-2t})+1}m_{1}(\Phi_{t})$. Similarly, we re-write the second term as
@@
$\leq 2m_{1}(\Phi_{t})\frac{(L-1)e^{-2t}}{(L-1)(1-e^{-2t})+1}+2\|\nabla\log(p_{t}^{V})(0)\|m_{0}(\Phi_{t})\,.$
@@
Using Lemma 21 and Proposition 22, we have that $\max(\Phi_{t})\leq\max(\Phi_{0})=\frac{(Z_{V})^{2}}{Z_{2V}}\max(\frac{\mathrm{d}\mu}{\mathrm{d}x})\leq\frac{(Z_{V})^{2}}{Z_{2V}}(\frac{L}{2\pi})^{d/2}$. Hence, if we pick $R=\left(Z_{2V}m_{0}(\Phi_{t})\Gamma(d/2)/Z_{V}^{2}\right)^{1/d}L^{-1/2}2^{1/2-1/d}$, we get as an upper-bound for $B$:
$B$ $\leq\frac{Z_{V}^{2}}{Z_{2V}}d\sqrt{L}2^{1/d-1/2}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\Gamma(d/2)^{-1/d}$
$\leq\frac{2Z_{V}^{2}}{Z_{2V}}\sqrt{dL}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\,.$
@@
$\partial m_{1}(\Phi_{t})\leq(d+1)\left(1+2\frac{(L-1)e^{-2t}}{(L-1)(1-e^{-2t})+1}\right)m_{1}(\Phi_{t})+2\|\nabla\log(p_{t}^{V})(0)\|m_{0}(\Phi_{t})+\frac{2Z_{V}^{2}}{Z_{2V}}\sqrt{Ld}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\,,$

### v3

@@ Appendix C Proof of Proposition 4
Appendix D Proof of Proposition  6 and Lemma 8
Before starting the proofs, we recall the following usefull lemma that bounds the second order moment of dissipative distributions.
Lemma 13
Let $V$ be such that $\langle\nabla V(x),x\rangle\geq a\|x\|^{2}-b$. Then for $\mu\propto e^{-V}$, denoting $m_{2}$ the second moment of $\mu$, it holds that $m_{2}\leq\frac{b+d}{a}$.
Proof.
Define the Laplacian of $\mu$ as $L(f)=\Delta f-\langle\nabla V,\nabla f\rangle$ for $f$ sufficiently smooth. By integration by parts, it holds that
$\int L(f)(x)\mathrm{d}\mu(x)=0\,.$
In particular, for $f(x)=\|x\|^{2}$, we recover $\int\langle\nabla V(x),\nabla f(x)\rangle\mathrm{d}\mu(x)=2d$. For $V$ dissipative, it implies
$a\int\|x\|^{2}\mathrm{d}\mu(x)\leq 2d+b\,$,
or equivalently $m_{2}\leq(b+2d)/a$. ∎
D.1 Proof of Lemma 8
Proof.
The first inequality was shown in the Lemma above. For the Fisher information, it holds that
$\mathcal{I}(\mu,\pi)$ $=\int\|\nabla V(x)-x\|^{2}\mathrm{d}\mu(x)\,$,
$\leq 2\int\langle\nabla V(x),\nabla V(x)e^{-V(x)}\rangle/Z_{V}\mathrm{d}x+2m_{2}\,$,
$=2\int\Delta V\mathrm{d}\mu(x)+2m_{2}\,$,
$\leq 2\beta d+2m_{2}\,.$
There remains to lower-bound $\mu(0)$. Denote $x^{*}$ a global minimizer of $V$. By dissipativity, it must hold that $\|x^{*}\|^{2}\leq b/a$. Now observe that
$V(x^{*})-V(0)$ $=\int_{0}^{1}(x^{*})^{\top}\nabla^{2}V(tx^{*})x^{*}\mathrm{d}t\,\leq\beta\|x^{*}\|^{2}\,.$
Combined with the fact that for all $x\in\mathbb{R}^{d}$, it holds that $V(x)-V(x^{*})\geq 0$, be recover that $V(x)-V(0)\geq-\beta\|x^{*}\|^{2}\geq-\beta b/a$. Furthermore, for $0<\delta<1$, we have
$V(x)-V(0)$ $=\int_{0}^{1}\langle\nabla V(tx),x\rangle\mathrm{d}t\,$,
$=\int_{\delta}^{1}\langle\nabla V(tx),x\rangle\mathrm{d}t+V(\delta x)-V(0)\,$,
$\geq\int_{\delta}^{1}\langle\nabla V(tx),tx\rangle/t\mathrm{d}t-\beta b/a\,$,
$\geq\int_{\delta}^{1}at\|x\|^{2}-b/t\mathrm{d}t-\beta b/a\,$,
$=a\|x\|^{2}/2-a\delta^{2}\|x\|^{2}/2+b\log(\delta)-\beta b/a\,.$
In particular, for $\delta=1/\sqrt{2}$, we obtain
$V(x)-V(0)\geq a\|x\|^{2}/4-b(\beta/a+\log(2)/2)\,.$
Hence, we have
$\mu(0)$ $=\frac{e^{-V(0)}}{\int e^{-V(x)}\mathrm{d}x}\,$,
$=\frac{1}{\int e^{-(V(x)-V(0))}\mathrm{d}x}\,$,
$\leq\frac{1}{\int e^{-a\|x\|^{2}/4+b(\beta/a+\log(2)/2)}\mathrm{d}x}$
$=e^{-b(\beta/a+\log(2)/2)}(a/2)^{d/2}(2\pi)^{-d/2}\,.$
Since $\beta/a\geq 1$ and $\log(2)/2\leq 1$, we recover that $\log(\mu(0)^{-2/d})\leq 4\beta b/ad+2\pi+\log(2/a)$.
D.2 Proof of Proposition  6
Recall that the intermediate scores read
$\nabla\log(p_{t}^{V})(z)=\frac{z-e^{-t}\mathbb{E}_{q_{t,z}}[Y]}{1-e^{-2t}}\,$,
with $q_{t,z}(x)\propto e^{-V(x)}e^{-\frac{\|e^{-t}x-z\|^{2}}{2(1-e^{-2t})}}$. In particular, if $V$ is dissipative with constant $a,b$ then it holds that
$\langle-\nabla\log(q_{t,z})(x),x\rangle$ $=\langle\nabla V(x)+\frac{e^{-t}(e^{-t}x-z)}{1-e^{-2t}},x\rangle\,$,
$\geq a\|x\|^{2}-b+\frac{e^{-t}}{1-e^{-2t}}(e^{-t}\|x\|^{2}-\langle z,x\rangle).$
Now recall that $e^{-t}\|x\|^{2}-\langle z,x\rangle\geq-e^{t}\|z\|^{2}/4$ which yields
$\langle-\nabla\log(q_{t,z})(x),x\rangle\geq a\|x\|^{2}-(b+\|z\|^{2}/(4(1-e^{-2t})))\,.$
Hence, using Lemma 13, it holds that
$\mathbb{E}_{q_{t,z}}[\|Y\|^{2}]\leq\frac{2b+\|z\|^{2}/(2(1-e^{-2t}))+d}{a}\,.$
Hence we recover that
$\|\nabla\log(p_{t}^{V})(z)\|^{2}$ $\leq\frac{2\|z\|^{2}}{(1-e^{-2t})^{2}}+\frac{2e^{-2t}\mathbb{E}_{q_{t,z}}[\|Y\|^{2}]}{(1-e^{-2t})^{2}}\,$,
$\leq\frac{2\|z\|^{2}}{(1-e^{-2t})^{2}}+\frac{e^{-2t}}{(1-e^{-2t})^{2}}\biggl(\frac{2(2b+d)}{a}+\frac{\|z\|^{2}}{(1-e^{-2t})a}\biggr)\,$,
$=\frac{\|z\|^{2}}{(1-e^{-2t})^{2}}\biggl(2+\frac{e^{-2t}}{a(1-e^{-2t})}\biggr)+\frac{2e^{-2t}(2b+d)}{a(1-e^{-2t})^{2}}\,.$
Similarly, recall that
$\nabla^{2}\log(p_{t}^{V})(z)-\nabla^{2}\log(\pi)(z)=\frac{e^{-2t}}{1-e^{-2t}}\left(\frac{\text{Cov}_{q_{t,z}}(X)}{1-e^{-2t}}-I_{d}\right)\,$,
from which we can deduce $\Delta\log(p_{t}^{2V})\leq\frac{e^{-2t}}{(1-e^{-2t})^{2}}\mathbb{E}_{q_{t,z}}[\|Y\|^{2}]$ which yields again
$\Delta\log(p_{t}^{2V})\leq\frac{e^{-2t}}{(1-e^{-2t})^{2}}\biggl(\frac{\|z\|^{2}}{2a(1-e^{-2t})}+\frac{2b+d}{a}\biggr)\,.$
Appendix E Proof of Lemma 7
Before starting the proof, we recall the result of Mikulincer and Shenfeld [2023].
Proposition 14
Let $\mu$ be a $\beta$-semi-log-convex probability distribution. Then, denoting $p_{t}^{V}$ the distribution of the forward process in Eq. 1 and $\pi$ the density of the standard Gaussian, it holds for all $z\in\mathbb{R}^{d}$ that
$\nabla^{2}\log(\pi)(z)-\nabla^{2}\log(p_{t}^{V})(z)\preceq\frac{(\beta-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}I_{d}\,.$ (28)
@@ Appendix E Proof of Lemma 7
In particular, we remark that $\nabla\log(Q_{t}(f))=\nabla\log(p_{t}^{V})-\nabla\log(\pi)$. Now, the quantity $\nabla\log(Q_{t}(f))$ was studied in Mikulincer and Shenfeld [2023] and they prove in Lemma 5 that for all $z$
$\nabla^{2}\log(Q_{t}(f))(z)\succeq\frac{(1-\beta)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}I_{d}\,$,
@@
$\nabla^{2}\log(\pi)(z)-\nabla^{2}\log(p_{t}^{V})(z)\preceq\frac{(\beta-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}I_{d}\,.$
@@
$\partial_{t}\Phi_{t}=\Phi_{t}\left(\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\right)\,.$
@@
$\partial_{t}\log(p_{t}^{V})=\Delta\log(p_{t}^{V})+\|\nabla\log(p_{t}^{V})\|^{2}-\langle\nabla\log(p_{t}^{V}),\nabla\log(\pi)\rangle-\Delta\log(\pi)\,.$
@@
$\partial_{t}\log(\Phi_{t})=\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(p_{t}^{2V})\|^{2}-\|\nabla\log(p_{t}^{V})\|^{2}\,.$
@@
$\partial_{t}\log(\Phi_{t})=\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\,.$
@@
$\partial_{t}m_{0}(\Phi_{t})$ $=\int\partial_{t}\Phi_{t}\mathrm{d}z$
@@
where we used Lemma 15 to compute $\partial_{t}\Phi_{t}$. Using integration by parts, the first term reads
@@
$\partial_{t}m_{0}(\Phi_{t})=2\int\langle\nabla\log(p_{t}^{V}),\nabla\Phi_{t}\rangle\mathrm{d}z-\int\langle\nabla\log(\pi),\nabla\Phi_{t}\rangle\mathrm{d}z\,.$
@@
$\partial_{t}m_{0}(\Phi_{t})$ $=\int\Delta\log(\pi)\Phi_{t}\mathrm{d}z-2\int\Delta\log(p_{t}^{V})\Phi_{t}\mathrm{d}z$
@@
Using Proposition 14, since $\mu$ is $\beta$-semi-log-convex, the term $(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))$ can be upper-bounded uniformly by $\frac{d(\beta-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}$ so we eventually get
$\partial_{t}m_{0}(\Phi_{t})\leq dm_{0}(\Phi_{t})\left(1+\frac{2(\beta-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}\right)\,.$
Hence we can use Gronwall’s lemma which yields
$m_{0}(\Phi_{t})\leq m_{0}(\Phi_{0})\exp\left(d\int_{0}^{t}\left(1+\frac{2(\beta-1)e^{-2s}}{(1-e^{-2s})(\beta-1)+1}\right)\mathrm{d}s\right)\,.$
@@
Finally, let us compute the integral above. Making the change of variable $u=e^{-2s}(\beta-1)$ we have $\mathrm{d}u=-2(\beta-1)e^{-2s}\mathrm{d}s$ which yields
$\int_{0}^{t}\frac{2(\beta-1)e^{-2s}}{(1-e^{-2s})(\beta-1)+1}\mathrm{d}s$ $=-\int_{\beta-1}^{(\beta-1)e^{-2t}}\frac{1}{\beta-u}\mathrm{d}u$
$=\left[\log(\beta-u)\right]_{\beta-1}^{(\beta-1)e^{-2t}}$
$=\log(\beta-(\beta-1)e^{-2t})$
$=\log(\beta(1-e^{-2t})+e^{-2t})\,.$
@@
$m_{0}(\Phi_{t})\leq\frac{e^{td}(Z_{V})^{2}}{Z_{2V}}(\beta(1-e^{-2t})+e^{-2t})^{d}\,.\qed$
@@
Let $z_{t}$ be a point where $\Phi_{t}$ attains its maximum and denote $M_{t}=\log(\Phi_{t})(z_{t})$. By the implicit function theorem, $z_{t}$ is differentiable hence we can compute $\partial_{t}M_{t}$ as
$\partial_{t}M_{t}$ $=\partial_{t}\log(\Phi_{t})(z_{t})+\langle\partial_{t}z_{t},\nabla\log(\Phi_{t})(z_{t})\rangle$
@@
Since $z_{t}$ is a maximum, we have in particular $\Delta\log\Phi_{t}(z_{t})\leq 0$ which implies that $M_{t}$ decreases. ∎
@@
If $\mu$ is $\beta$-semi-log-convex then it holds that $\frac{\mathrm{d}\mu}{\mathrm{d}z}\leq\left(\frac{\beta}{2\pi}\right)^{\frac{d}{2}}$.
@@
where $V_{*}$ the minimum of $V$ attained for some $z_{*}$. By definition $e^{-(V(z)-V_{*})}\leq 1$ for all $z$. Furthermore, since $V$ verifies $\nabla^{2}V\preceq\beta I_{d}$, we are ensured that
$V(z)-V_{*}\leq\beta\frac{\|z-z_{*}\|^{2}}{2}\,$,
@@
$\frac{1}{\int_{z}e^{-(V(z)-V_{*})}\mathrm{d}z}\leq\left(\frac{\beta}{2\pi}\right)^{\frac{d}{2}}\,.\qed$
Using the previous result, we can derive an upper-bound on the integrated squared gradient at 0.
@@
Let $\mu\propto e^{-V}$ be a $\beta$-semi-log-convex measure. Denoting $\mu(0)$ the density of $\mu$ with respect to the Lebesgue measure at 0, it holds that
$\int_{0}^{t}\|\nabla\log(p_{s}^{V})\|^{2}(0)\mathrm{d}s\leq-\log\left(\mu(0)\right)+\frac{d}{2}\log\left(\frac{\beta}{2\pi}\right)\,.$
@@
$\partial_{t}p_{t}^{V}=\nabla\cdot\left(p_{t}^{V}\nabla\log\left(\frac{p_{t}^{V}}{\pi}\right)\right)\,$,
@@
$\partial_{t}\log(p_{t}^{V})$ $=\frac{\Delta p_{t}^{V}}{p_{t}^{V}}-\langle\nabla\log(p_{t}^{V}),\nabla\log(\pi)\rangle-\Delta\log(\pi)$
@@
$\partial_{t}\log(p_{t}^{V})(0)=\Delta\log(p_{t}^{V})(0)+\|\nabla\log(p_{t}^{V})\|^{2}(0)-\Delta\log(\pi)(0)\,$,
@@
Using the uniform upper-bound of Proposition 14, the second term can upper-bounded as
$\int_{0}^{t}\Delta\log(\pi)(0)-\Delta\log(p_{t}^{V})(0)\mathrm{d}s$ $\leq d\int_{0}^{t}\frac{(\beta-1)e^{-2s}}{(1-e^{-2s})(\beta-1)+1}\mathrm{d}s$
$=\frac{d}{2}(\beta-1)\int_{e^{-2t}}^{1}\frac{1}{\beta-u(\beta-1)}\mathrm{d}u$
$=\frac{d}{2}\log(\beta(1-e^{-2t})+e^{-2t})\,.$
Furthermore, Proposition 14 shows that $\nabla^{2}\log(p_{t}^{V})\preceq\frac{\beta}{\beta(1-e^{-2t})+e^{-2t}}$. Thus, using Proposition 17, we recover that $\log(p_{t}^{V})(0)\leq\frac{d}{2}\log\left(\frac{\beta}{\beta(1-e^{-2t})+e^{-2t}}\right)-\frac{d}{2}\log(2\pi)$. In particular, we recover
$\int_{0}^{t}\|\nabla\log(p_{s}^{V})\|^{2}(0)\mathrm{d}s\leq-\log\left(\frac{\mathrm{d}\mu}{\mathrm{d}x}(0)\right)+\frac{d}{2}\log(\beta)-\frac{d}{2}\log(2\pi)\,.\qed$
@@
Let $\mu$ be a $\beta$-semi-log-convex measure with finite second moment $m_{2}$. It holds that
$m_{1}(\Phi_{t})\leq\frac{(Z_{V})^{2}}{Z_{2V}}e^{t(d+1)}(\beta(1-e^{-2t})+e^{-2t})^{d}\left(\sqrt{m_{2}}+\sqrt{-2\log\left(\mu(0)\right)+d\log\left(\frac{\beta}{2\pi}\right)}+2\sqrt{d\beta}\right)\,.$
@@
$\partial_{t}m_{1}(\Phi_{t})=\int\Phi_{t}\left(\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\right)\|z\|\mathrm{d}z\,.$
@@
$\partial_{t}m_{1}(\Phi_{t})=\int\langle 2\nabla\log(p_{t}^{V})-\nabla\log(\pi),\nabla\Phi_{t}\rangle\|z\|\mathrm{d}z-\int\langle\nabla\Phi_{t},\frac{z}{\|z\|}\rangle\mathrm{d}z\,.$
@@
Using the upper bound given in Proposition 14, we get $2\int\Phi_{t}(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|\mathrm{d}z\leq 2d\frac{(\beta-1)e^{-2t}}{(\beta-1)(1-e^{-2t})+1}m_{1}(\Phi_{t})$. Similarly, we re-write the second term as
@@
$\leq 2m_{1}(\Phi_{t})\frac{(\beta-1)e^{-2t}}{(\beta-1)(1-e^{-2t})+1}+2\|\nabla\log(p_{t}^{V})(0)\|m_{0}(\Phi_{t})\,.$
@@
Using Lemma 16 and Proposition 17, we have that $\max(\Phi_{t})\leq\max(\Phi_{0})=\frac{(Z_{V})^{2}}{Z_{2V}}\max(\frac{\mathrm{d}\mu}{\mathrm{d}x})\leq\frac{(Z_{V})^{2}}{Z_{2V}}(\frac{\beta}{2\pi})^{d/2}$. Hence, if we pick $R=\left(Z_{2V}m_{0}(\Phi_{t})\Gamma(d/2)/Z_{V}^{2}\right)^{1/d}\beta^{-1/2}2^{1/2-1/d}$, we get as an upper-bound for $B$:
$B$ $\leq\frac{Z_{V}^{2}}{Z_{2V}}d\sqrt{\beta}2^{1/d-1/2}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\Gamma(d/2)^{-1/d}$
$\leq\frac{2Z_{V}^{2}}{Z_{2V}}\sqrt{d\beta}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\,.$
@@
$\partial_{t}m_{1}(\Phi_{t})\leq(d+1)\left(1+2\frac{(\beta-1)e^{-2t}}{(\beta-1)(1-e^{-2t})+1}\right)m_{1}(\Phi_{t})\\ +2\|\nabla\log(p_{t}^{V})(0)\|m_{0}(\Phi_{t})+\frac{2Z_{V}^{2}}{Z_{2V}}\sqrt{\beta d}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\,$, (29)